OpenStudy (anonymous):

Helllllllpppp, MATH, PLEASE The area of a circle is 28in squared. The length of the radius is _____in. 3 4.2 6.4 8.9

OpenStudy (anonymous):

(A)rea = (pi)(r^2) = 28 r = sqrt(28/pi) = sqrt(28/3.14) = 2.98 which is about 3

OpenStudy (anonymous):

thank you soo much, I'm terrible with circles & all the formulas, would you wanna help me with another one.

OpenStudy (anonymous):

uw! Sure!

OpenStudy (anonymous):

What is the lateral area of a cylinder with a radius of 5 ft and a height of 8 ft? 157ft2 251.2 ft2 96ft2 408.2 ft2

OpenStudy (anonymous):

The area of the "sides" is 2(pi)5 times 8 = 80 * pi = 251.32741228718345907701147066236 The area of the "top" and "bottom" is: 2 x (pi)r^2 = 50(pi) = 157.07963267948966192313216916398

OpenStudy (anonymous):

The lateral area is just the first one, 251.33. It's "without top and bottom". It's just the "label" part if you think of a can of soup.

OpenStudy (anonymous):

The "whole area" is 408, the "added up" amount.